OpenStudy (vera_ewing):

Which of the following is the conjugate of a complex number with 2 as the real part and 7 as the imaginary part? (5 points) A. 2 + 7i B. -2 - 7i C. -2 + 7i D. 2 - 7i

OpenStudy (vera_ewing):

@paki PLease help!

OpenStudy (vera_ewing):

@TheSmartOne help!

OpenStudy (vera_ewing):

@ganeshie8 please help me!

OpenStudy (irishboy123):

z = 2 + 7i z* = 2 - 7i

OpenStudy (vera_ewing):

@confluxepic please help me!

OpenStudy (irishboy123):

z* is the conjugate if z = 2 + 7i